OpenStudy (anonymous):

Prove: 1/1+sinx+1/1-sinx=2sec^2x 1/1+sinx+1/1-sinx=...?

OpenStudy (luigi0210):

I'm confused.. what's your equation?

Parth (parthkohli):

\[\dfrac{1}{1 + \sin(x)} + \dfrac{1}{1 - \sin(x)} = 2 \sec^2(x)\]

Parth (parthkohli):

Try to get a common denominator on the left-hand side. Also try to pronounce the right-hand side loud. lol

OpenStudy (luigi0210):

\[\frac{ 1 }{ sinx+1 }+\frac{ 1 }{ 1-sinx }=2\sec^2x\]

OpenStudy (luigi0210):

I've done this proof so many times >.>

Parth (parthkohli):

haha yes

Parth (parthkohli):

\[\dfrac{1 - \sin(x) }{(1 + \sin(x))(1 - \sin(x))} + \dfrac{1 + \sin(x)}{(1 - \sin(x))(1 + \sin(x))}\]

Parth (parthkohli):

You'd get something special in the denominator which is related to \(\sec^2(x)\).

OpenStudy (luigi0210):

and a 2

Parth (parthkohli):

\[\dfrac{1 - \sin(x) + 1 + \sin(x)}{\text{I'd leave this part to you}}\]

OpenStudy (jhannybean):

SO THATS HOW YOU WRITE TEXT IN THERE omg.

Parth (parthkohli):

@Luigi0210 Yeah, the 2 comes from the numerator.

Parth (parthkohli):

Trig is so interesting :')

Parth (parthkohli):

@isuckatmath9999

OpenStudy (luigi0210):

Trig loves you guys but hates me

OpenStudy (anonymous):

im thinking -- err processing

Parth (parthkohli):

Take your time.

OpenStudy (anonymous):

my interesting version of math left a denominator of 1... lol

Parth (parthkohli):

Naw

Parth (parthkohli):

The denominator is \((1 - \sin(x))(1 + \sin(x))\). You can simplify that so easily.

OpenStudy (anonymous):

normal people can simplify that easily lol

Parth (parthkohli):

\[(x - y)(x + y) = x^2 - y^2 \]

Parth (parthkohli):

Makes sense?

OpenStudy (anonymous):

yes

Parth (parthkohli):

So what would \((1 - \sin(x))(1 + \sin(x))\) be?

OpenStudy (anonymous):

1^2-sin(x)^2 1-sin(x)^2 ?

Parth (parthkohli):

Indeed, indeed. See any identity there?

Parth (parthkohli):

\[? = 1 - \sin^2(x)\]

OpenStudy (anonymous):

uhhmm

Parth (parthkohli):

Recall!

OpenStudy (anonymous):

cosx ?

OpenStudy (anonymous):

wawait that would need a sqrt

Parth (parthkohli):

Yeah, close.

OpenStudy (anonymous):

coverse sine?

Parth (parthkohli):

Why are you complicating it?\[\sin^2(x) + \cos^2(x) = 1\]

OpenStudy (jhannybean):

xD

OpenStudy (anonymous):

because I get easily confused?

Parth (parthkohli):

Err man.

OpenStudy (anonymous):

and i still hav eno idea why its 2sec^2x =/

Parth (parthkohli):

\[\sin^2(x) + \cos^2(x) = 1 \Rightarrow \cos^2(x) = 1 - \sin^2(x)\]

Parth (parthkohli):

You have\[\dfrac{1 + \sin(x) + 1 - \sin(x)}{1 - \sin^2(x)}\]Right?

OpenStudy (anonymous):

right

Parth (parthkohli):

Can you simplify the numerator?

OpenStudy (anonymous):

I'm not sure how. I don't have a chart of all of these things.

Parth (parthkohli):

\[1 + \sin(x) + 1 - \sin(x)\]!!!

OpenStudy (anonymous):

2

Parth (parthkohli):

Good. And the denominator?

OpenStudy (anonymous):

1-sin^2x?

Parth (parthkohli):

Yeah, but I told you that \(1 - \sin^2(x) = \cos^2(x)\)

OpenStudy (anonymous):

why do i need it to = cos^2x?

Parth (parthkohli):

Because \(\sec^2(x)\) is related to \(\cos^2(x)\), RIGHT?

OpenStudy (anonymous):

right?

Parth (parthkohli):

\[\dfrac{2}{\cos^2(x)}\]is what you have

Parth (parthkohli):

Now recall that\[\dfrac{1}{\cos^2(x)} = \sec^2(x)\]

OpenStudy (anonymous):

ooohhhh

OpenStudy (anonymous):

I get it!!! :D